# Planting and MCMC Sampling from the Potts model

### v1

@@ 4. Proof for WSM in the disordered regime
To show WSM in the disordered regime, we follow the strategy outlined in Section 1.2. We begin with some relevant definitions. Let $G=(V,E)$ be a graph. Recall, that for a vertex $v$ and $\ell>0$, $B_{\ell}(v)$ denotes the $-\ell$ ball around $v$ in $G$. Also, denote by $S_{\ell}(v)$ the set of vertices at distance exactly $\ell$ from $v$. For a set of vertices $S$, we let $E(S)$ the set of edges with both endpoints in $S$.
@@
For a boundary condition $\eta$ induced by $F\setminus B_{\ell}(v)$, a nontrivial boundary component of $\eta$ is a component of $(V,F\backslash E(B_{\ell}(v)))$ containing at least two vertices of $S_{\ell}(v)$.
@@
For a graph $G\in\mathcal{G}_{n,d}$ and $\sigma\in[q]^{n}$, let $\boldsymbol{F}(G,\sigma)$ denote the random subset of edges obtained by keeping each monochromatic edge of $G$ under $\sigma$ with probability $p$. Fix $\epsilon>0$ and let $K=\lceil 10/\epsilon\rceil$. We will consider the disordered planted graph $\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}$ and show that for an arbitrary vertex $v$ it holds that
@@
It remains to prove (4.1). By Lemma 2.3, we have $\big{\|}\nu^{\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\nu_{\mathrm{d}}\big{\|}_{1}+\big{\|}\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\rho_{\mathrm{d}}\big{\|}_{1}=o(1)$ w.p. $1-\mathrm{e}^{-\Omega(n)}$. Consider arbitrary $\nu$ and $\rho$ with $\big{\|}\nu-\nu_{\mathrm{d}}\big{\|}_{1}+\big{\|}\rho-\rho_{\mathrm{d}}\big{\|}_{1}=o(1)$. Conditioned on $\nu^{\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}=\nu$ and $\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}=\rho$, we have that $(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}})$ is uniformly random among all pairs $(G,\sigma)$ with $\nu^{\sigma}=\nu$ and $\rho^{G,\sigma}=\rho$.
To proceed, fix $\sigma$ with $\nu^{\sigma}=\nu$ and condition on $\hat{\boldsymbol{\sigma}}_{\mathrm{d}}=\sigma$. Condition further that $\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}=\rho$ and on the induced graph on $B_{\ell}(v)$. For convenience, let $\tilde{\boldsymbol{F}}$ denote the (random) set of edges outside $B_{\ell}(v)$ after the percolation step, i.e., the set $\boldsymbol{F}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}\big{)}\backslash E(B_{\ell}(v))$.
Let $T:=\lfloor n^{1/2-\epsilon}\rfloor$. For $t=1,\ldots,T$, we explore the (union of the) components of vertices $u\in S_{\ell}(v)$ in the graph $(V,\tilde{\boldsymbol{F}})$ by revealing at each step an edge incident to a vertex belonging to any of these (chosen arbitrarily). We denote by $A_{t}$ the set of “active vertices” in these components that still have unmatched half-edges, so that $|A_{t}|=0$ implies that all the components have been fully explored. Let $A_{\leq t}=\cup_{t^{\prime}\leq t}A_{t}$ denote the set of all vertices that have been active up to time $t$. The process is described precisely in Figure 1.
{mdframed}
Exploration of $(V,\tilde{\boldsymbol{F}})$ starting from $S_{\ell}(v)$ Let $t=0$, $T=\lfloor n^{1/2-\epsilon}\rfloor$, and $A_{0}=S_{\ell}(v)$.While $|A_{t}|\neq 0$ or $t\leq T$:
  • If $|A_{t}|=0$, pick any unmatched half-edge $(w,r)$ and run Match$(w,r)$;
    increase $t$ by 1 and set $A_{t}=\emptyset$
  • Otherwise, pick $w\in A_{t}$. For $r=1,\ldots,d$:
    If the half-edge $(w,r)$ is not matched:
    increase $t$ by 1;
    match $(w,r)$ to another half-edge, let $z=\textsc{Match}$$(w,r)$.
    if $\sigma(z)=\sigma(w)$, w.p. $p=1-\mathrm{e}^{-\beta}$ set $A_{t}\leftarrow A_{t-1}\cup\{z\}$; else set $A_{t}=A_{t-1}$.
    Remove $w$ and any other vertices from $A_{t}$ which don’t have unmatched half-edges.
Match$(w,r)$
@@
  • Let $i=\sigma(w)$. Choose a colour $j\in[q]$ from the distribution $\{\frac{M_{ij,t}}{\sum_{j^{\prime}}M_{ij^{\prime},t}}\}_{j\in[q]}$.
  • From the set of unmatched half-edges from colour $j$ to $i$, choose one u.a.r., say $(z,r^{\prime})$.
  • match $(z,r^{\prime})$ with $(w,r)$, and return the vertex $z$.
Figure 1.
Note that at any time $t$, for $i,j\in[q]$, there is a number $M_{ij,t}$ of prescribed half-edges with one endpoint of colour $i$ that needs to be matched to a vertex of colour $j$ (note, this goes down when we reveal such an edge connecting colours $i$ and $j$). The revealing of an edge at time $t$ adjacent to an active vertex $w$ with colour $\sigma(w)=i$ has therefore two stages: we first reveal the other endpoint $z$ in the graph $\hat{\boldsymbol{G}}(\sigma)$, chosen proportionally to the counts $M_{ij,t}$; if $\sigma(w)=\sigma(z)$, we further toss a coin with probability $p=1-\mathrm{e}^{-\beta}$ to determine whether the edge belongs to $\tilde{\boldsymbol{F}}$, and add $z$ to the set of active vertices accordingly. The key point here is that we only need to reveal the components involving $S_{\ell}(v)$ in the graph $(V,\tilde{\boldsymbol{F}})$, rather than the components in $\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}})$ (which are much bigger). Indeed, as we will see shortly the former is dominated by a subcritical branching process and therefore at most $O(\log n)$ vertices are revealed for each of the vertices in $S_{\ell}(v)$.
@@
(To see how the $o(1)$ term arises, see the proof of Lemma B.1.) It follows that the sum $\sum^{T}_{t=1}\boldsymbol{1}_{t}$ is dominated above by a sum of $T$ independent coin tosses with probability $\frac{1}{n^{1/2+\epsilon/2}}$, whereas the sum $\sum^{T}_{t=1}\boldsymbol{1}_{t}^{\prime}$ is dominated above by a sum of $T$ independent coin tosses with probability $\frac{\mathrm{e}^{\beta}-1}{\mathrm{e}^{\beta}+q-1}+o(1)<\frac{1-2\kappa}{d-1}$ for some small constant $\kappa=\kappa(q,d,\beta)>0$ (using that $\mathrm{e}^{\beta}-1<q/(d-2)$ for $\beta<\beta_{u}^{\prime}$). Therefore, recalling that $K=\lceil 10/\epsilon\rceil$ and $T=\lfloor n^{1/2-\epsilon}\rfloor$, we have
@@
Note that the event $|A_{T}|>0$ implies that $|A_{t}|>0$ for all $t=0,1,\ldots,T$ which in turn implies that at least one vertex gets removed from the set of active vertices every (at most) $d-1$ time steps; indeed, when exploring the neighbours of a vertex $w\in A_{t}$ at a particular time $t$, there are at most $d-1$ half-edges incident to $w$ remaining unmatched since one of the $d$ half-edges of $w$ has been previously matched (to activate $w$). We therefore have the inequality $0\leq|A_{T}|\leq|S_{\ell}(v)|+\sum^{T}_{t=1}\boldsymbol{1}_{t}^{\prime}-\frac{T-(d-1)}{d-1}$. Since $|S_{\ell}(v)|\leq d(d-1)^{\ell}\leq dn^{1/2-2\epsilon}\leq\frac{\kappa T}{d-1}$, it follows that
@@
Combining these, we obtain that with probability $\geq 1-2/n^{5}$ both of $|A_{T}|=0$ and $\sum^{T}_{t=1}\boldsymbol{1}_{t}<K$ hold. The first implies that by time $T$ all the components of $S_{\ell}(v)$ in the graph $(V,\tilde{\boldsymbol{F}})$ have been explored, and the second implies that all but (at most) $K$ of these vertices belong to distinct components. This finishes the proof of (4.1) and hence completes the proof of the lemma. ∎
We next use an argument of Blanca and Gheissari [BG] which shows a correlation decay bound for $K$-shattered configurations in a treelike graph in terms of the existence of two root-to-leaf paths. The argument in [BG] is technically proved for $\beta<\beta_{u}$ using some uniqueness estimates, but it can be extended it to the regime $\beta\leq\beta_{c}$ by observing that in a treelike ball, a $K$-shattered configuration induces a distribution that is comparable to percolation on the tree with parameter $\hat{p}$ (this coincides with the cut-edge update probability in the definition of the random cluster dynamics). A key feature of the correlation decay bound in [BG] is the existence of two paths which gives the $\hat{p}^{2\ell}$ term in the lemma below, and which ultimately allows us to make the probability less than $\frac{1}{100m}$ that is required for WSM, while keeping the distance $\ell$ less than $\frac{1}{2}\log_{d-1}n$ that is the threshold for having treelike neighbourhoods. We give the proof of the following lemma in Appendix C.
@@
Lemma 4.3 gives that, whp over $G\sim\mathcal{G}_{n,d}$, $\big{|}\pi_{G,v,\ell}^{-}(1_{e})-\pi_{G,\mathrm{d}}(1_{e})\big{|}\leq C\hat{p}^{2l}+\frac{1}{n^{3}}$ for any edge $e$ and vertex $v$ incident to it. By the choice of $\delta,\epsilon$ and $\ell$, we get that $p^{2l}\leq n^{-(1+\epsilon/2)}/\hat{p}$, so for large enough $n$ we have $C{\hat{p}}n^{-(1+\epsilon/2)}+n^{-3}\leq\frac{1}{100m}$, concluding the proof of WSM within the disordered phase. ∎
@@ 5. Proof for WSM within the ordered regime
For a graph $G$, let ${\mathcal{C}}_{1}(G)$ denote the largest component of $G$ (breaking ties arbitrarily), and more generally, ${\mathcal{C}}_{i}(G)$ denotes the $i$-th largest component of $G$. It is a standard fact [perc1, perc2, coja2023] that the size of the largest component in a percolated random $d$-regular graph is controlled by the extinction probability $\varphi$ of a Galton-Watson process with distribution $\mathrm{Bin}(d-1,p)$ on the tree. For $p\in(\tfrac{1}{d-1},1)$, define $\varphi=\varphi(p),\chi=\chi(p)$ and $\hat{\varphi}=\hat{\varphi}(p)$ be the unique solutions on the interval $(0,1)$ to the following
@@
In particular, the size of the largest component is roughly equal to $\chi n+o(n)$; see [perc1, perc2] for more details. We also remark that the quantity $\hat{\varphi}$ is the extinction probability of a slightly modified GW-tree where every vertex has degree $d$ apart from the root which has degree $d-2$; this will be relevant later for our path considerations.
In the lemma below, we show a similar result for the size of the giant component for slightly non-regular degree sequences $\boldsymbol{d}$, where we allow a small set $S$ of vertices to have degree less than $d$. A variant of percolation we will use is the exact edge model for random graphs with a given degree sequence $\boldsymbol{d}=(d_{1},\dots,d_{n})$ and edge count $m\leq\tfrac{1}{2}\sum d_{i}$. Let $\tilde{\boldsymbol{G}}_{n,\mathbf{d},m}$ be the random graph model where each $i\in[n]$ is associated to $d_{i}$ half-edges; then we choose a subset of $2m$ half-edges and a matching of these uniformly at random; the final (multi)graph is obtained by projecting the edges on the vertex set $[n]$. The proof of the lemma is deferred to Appendix D.
@@
    If $p<1/(d-1)$, $\mathbb{P}\big{[}{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{1/2}]\leq\mathrm{e}^{-\Omega(n)}$.
@@
    If $p>1/(d-1)$, $\mathbb{P}\big{[}{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq(\chi-\delta)n]\geq 1-\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}\big{[}{\mathcal{C}}_{2}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{2/3}]\leq\mathrm{e}^{-\Omega(n)}$. Moreover, with $\varphi=\varphi(p)$ as in (5.1), for any subset $S^{\prime}\subseteq S$,
@@
The following lemma was shown in [coja2023]; intuituively, and as we will see later in detail, it says that in the ordered regime one color class is in the supercritical regime, while the others in subcritical.
Recall that $\nu_{\mathrm{o}}$ and $\rho_{\mathrm{o}}$ are the vertex and edge statistics for the ordered phase respectively, defined in (2.1) and (2.5).
Lemma 5.2 ([coja2023, Lemma 5.6]).
@@
Let $G=(V,E)$ be a graph. For a subset of edges $F\subseteq E$ and a vertex $v$, we let $F_{v}$ be the edges in $F$ incident to $v$. We will also consider the largest component ${\mathcal{C}}$ of the graph $(V,F)$, i.e. the component with the largest number of vertices (breaking ties arbitrarily). We say that ${\mathcal{C}}$ avoids a path $P$ if there is no vertex $v\in{\mathcal{C}}$ that belongs to $P$. The following lemma will be used to show that, in the RC ordered phase, the largest component (which turns out to have size $\Omega(n)$) is “very close” to every vertex $v$.
@@
$\mathcal{A}_{v}=\Big{\{}F\subseteq E(G)\,\Big{|}\,\begin{array}[]{c}\mbox{there exists a path of length $\ell$ from $v$ in $G$ which}\\ \mbox{avoids the largest component of $(V\backslash\{v\},F\backslash F_{v})$}\end{array}\Big{\}}.$
Then, for the RC ordered phase of $G$, $\pi_{G,\mathrm{o}}\big{(}\bigcup_{v\in V(G)}\mathcal{A}_{v}\big{)}\leq nd(d-1)^{\ell-1}(A+\varepsilon)^{\ell}$ where $A=\max_{i\in[q]}\big{\{}1-(1-{\small\sqrt{\hat{\varphi}_{1}})}\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}\big{\}}$ with $\rho_{\mathrm{o}},\nu_{\mathrm{o}}$ as in (2.1) and $\hat{\varphi}_{1}=\hat{\varphi}(p_{1})$ as in (5.1) for $p_{1}=p\frac{\rho_{\mathrm{o},11}}{\nu_{\mathrm{o},1}}$.
To parse the bound on $\pi_{G,\mathrm{o}}\big{(}\bigcup_{v\in V(G)}\mathcal{A}_{v}\big{)}$ note that the factor $nd(d-1)^{\ell-1}$ is just an upper bound on the total number of paths of length $\ell$. The factor $A^{\ell}$ controls the probability that a particular path $P$ avoids the largest component. The term $\hat{\varphi}_{1}$ in $A$ is the probability that the Galton-Watson process dies out from a vertex $v$ of the path that belongs to the supercritical color class 1 (note that $v$ has “remaining” degree $d-2$ other than its neighbours on the path, and that the relevant percolation parameter is given by $p_{1}$). The term $\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}$ accounts for the frequency that we encounter vertices in the supercritical color class 1 along the path $P$, which follows from the definition of the planting model. The exact quantitative dependence of $A$ to $\hat{\varphi}_{1}$ and $\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}$ follows by a more technical calculation involving eigenvalues. Later, in Lemma 5.4, we will show that $A$ can be made less than $1/d^{5}$ by taking $q$ sufficiently large.
Proof of Lemma 5.3.
@@
From this, we obtain by Lemma 2.2 that $\mathbb{P}\big{[}\boldsymbol{F}\big{(}\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big{)}\in\bigcup_{v\in V}\mathcal{A}_{v}\mid\boldsymbol{G}\big{]}\leq n\ln(n)d\big{(}d-1)^{\ell-1}(A+\tfrac{\varepsilon}{2}\big{)}^{\ell}$; since $\boldsymbol{F}\big{(}\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}})$ has distribution $\hat{\pi}_{{\boldsymbol{G}},\mathrm{o}}$, Lemma 3.3 yields that, whp over $\boldsymbol{G}$, it holds that $\pi_{\boldsymbol{G},\mathrm{o}}(\bigcup_{v}\mathcal{A}_{v})\leq\pi_{\boldsymbol{G},\mathrm{o}}(\bigcup_{v}\mathcal{A}_{v})+\mathrm{e}^{-\Omega(n)}\leq nd(d-1)^{\ell-1}(A+\varepsilon)^{\ell}$.
@@
Let $P$ be an arbitrary sequence of vertices $(v_{0},v_{1},\ldots,v_{\ell})$, together with a $q$-colouring $\tau=(\tau_{0},\ldots,\tau_{\ell})$. Let also $\xi$ be a set of paired half-edges that turn $v_{0},\ldots,v_{\ell}$ into a path in that order. Denote by

### v2

@@ 4. Proof for WSM in the disordered regime
To show WSM in the disordered regime, we follow the strategy outlined in Section 1.2. We begin with some relevant definitions. Let $G=(V,E)$ be a graph. Recall, that for a vertex $v$ and $\ell>0$, $B_{\ell}(v)$ denotes the radius-$\ell$ ball around $v$ in $G$. Also, denote by $S_{\ell}(v)$ the set of vertices at distance exactly $\ell$ from $v$. For a set of vertices $S$, we let $E(S)$ the set of edges with both endpoints in $S$.
@@
For a graph $G\in\mathcal{G}_{n,d}$ and $\sigma\in[q]^{n}$, let $\boldsymbol{F}(G,\sigma)$ denote the random subset of edges obtained by keeping each monochromatic edge of $G$ under $\sigma$ with probability $p$. Fix $\epsilon>0$ and let $K=\lceil 10/\epsilon\rceil$. We will consider the disordered planted graph $(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}})$ and show that for an arbitrary vertex $v$ it holds that
@@
It remains to prove (4.1). By Lemma 2.3, we have $\big\|\nu^{\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\nu_{\mathrm{d}}\big\|_{1}+\big\|\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}-\rho_{\mathrm{d}}\big\|_{1}=O(n^{-1/3})$ w.p. $1-\mathrm{e}^{-\Omega(n^{1/3})}$. Consider arbitrary $\nu$ and $\rho$ with $\big\|\nu-\nu_{\mathrm{d}}\big\|_{1}+\big\|\rho-\rho_{\mathrm{d}}\big\|_{1}=O(n^{-1/3})$. Conditioned on $\nu^{\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}=\nu$ and $\rho^{\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}}=\rho$, we have that $(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}})$ is uniformly random among all pairs $(G,\sigma)$ with $\nu^{\sigma}=\nu$ and $\rho^{G,\sigma}=\rho$.
[…]
@@
Let $i=\sigma(w)$. Choose a colour $j\in[q]$ from the distribution $\{\frac{M_{ij,t}}{\sum_{j^{\prime}}M_{ij^{\prime},t}}\}_{j\in[q]}$. From the set of unmatched half-edges from colour $j$ to $i$, choose one u.a.r., say $(z,r^{\prime})$. match $(z,r^{\prime})$ with $(w,r)$, and return the vertex $z$.
Figure 1. The exploration process of the percolated planted graph used in the proof of Lemma 4.2. This is a combination of exposing first an edge in the planted configuration model, captured by the routine Match(,)($\cdot$$\cdot$), followed by a percolation step with probability $p=1-\mathrm{e}^{-\beta}$ whenever the edge is monochromatic.
[…]
@@
It follows that the sum $\sum^{T}_{t=1}\boldsymbol{1}_{t}$ is dominated above by a sum of $T$ independent coin tosses with probability $\frac{1}{n^{1/2+\epsilon/2}}$, whereas the sum $\sum^{T}_{t=1}\boldsymbol{1}_{t}^{\prime}$ is dominated above by a sum of $T$ independent coin tosses with probability $\frac{\mathrm{e}^{\beta}-1}{\mathrm{e}^{\beta}+q-1}+o(1)<\frac{1-2\kappa}{d-1}$ for some small constant $\kappa=\kappa(q,d,\beta)>0$ (using that $\mathrm{e}^{\beta}-1<q/(d-2)$ for $\beta<\beta_{u}^{\prime}$). Therefore, recalling that $K=\lceil 10/\epsilon\rceil$ and $T=\lfloor n^{1/2-\epsilon}\rfloor$, we have
@@
Note that the event $|A_{T}|>0$ implies that $|A_{t}|>0$ for all $t=0,1,\ldots,T$ which in turn implies that at least one vertex gets removed from the set of active vertices every (at most) d1d-1 time steps; indeed, when exploring the neighbours of a vertex $w\in A_{t}$ at a particular time $t$, there are at most d1d-1 half-edges incident to $w$ remaining unmatched since one of the $d$ half-edges of $w$ has been previously matched (to activate $w$). We therefore have the inequality $0\leq|A_{T}|\leq|S_{\ell}(v)|+\sum^{T}_{t=1}\boldsymbol{1}_{t}^{\prime}-\frac{T-(d-1)}{d-1}$. Since $|S_{\ell}(v)|\leq d(d-1)^{\ell}\leq dn^{1/2-2\epsilon}\leq\frac{\kappa T}{d-1}$, it follows that
@@
[…]
We next use an argument of Blanca and Gheissari [9] which shows a correlation decay bound for $K$-shattered configurations in a treelike graph in terms of the existence of two root-to-leaf paths. The argument in [9] is technically proved for $\beta<\beta_{u}$ using some uniqueness estimates, but it can be extended it to the regime $\beta\leq\beta_{c}$ by observing that in a treelike ball, a $K$-shattered configuration induces a distribution that is comparable to percolation on the tree with parameter $\hat{p}$ (this coincides with the cut-edge update probability in the definition of the random cluster dynamics). A key feature of the correlation decay bound in [9] is the existence of two paths which gives the $\hat{p}^{2\ell}$ term in the lemma below, and which ultimately allows us to make the probability less than $\frac{1}{100m}$ that is required for WSM, while keeping the distance $\ell$ less than $\frac{1}{2}\log_{d-1}n$ that is the threshold for having treelike neighbourhoods. We give the proof of the following lemma in Appendix C.
@@
Lemma 4.3 gives that, whp over $G\sim\mathcal{G}_{n,d}$, for any edge $e$ and vertex $v$ incident to it, it holds that $\big|\pi_{G,v,\ell}^{-}(1_{e})-\pi_{G,\mathrm{d}}(1_{e})\big|\leq C\hat{p}^{2l}+\frac{1}{n^{3}}$ . By the choice of $\delta,\epsilon$ and $\ell$, we get that $\hat{p}^{2l}\leq n^{-(1+\epsilon/2)}/(\hat{p}^{2})$, so for large enough $n$ we have $C{\hat{p}}^{-2}n^{-(1+\epsilon/2)}+n^{-3}\leq\frac{1}{100m}$, concluding the proof of WSM within the disordered phase. ∎
@@ 5. Proof for WSM within the ordered regime
For a graph $G$, let ${\mathcal{C}}_{1}(G)$ denote the largest component of $G$ (breaking ties arbitrarily), and more generally, ${\mathcal{C}}_{i}(G)$ denote the $i$-th largest component of $G$. It is a standard fact [3, 29, 14] that the size of the largest component in a percolated random $d$-regular graph is controlled by the extinction probability $\varphi$ of a Galton-Watson process with distribution $\mathrm{Bin}(d-1,p)$ on the tree. For $p\in(\tfrac{1}{d-1},1)$, define $\varphi=\varphi(p),\chi=\chi(p)$ and $\hat{\varphi}=\hat{\varphi}(p)$ to be the unique solutions on the interval $(0,1)$ to the following
@@
In particular, the size of the largest component is roughly equal to $\chi n+o(n)$; see [3, 29] for more details. We also remark that the quantity $\hat{\varphi}$ is the extinction probability of a slightly modified GW-tree where every vertex has degree $d$ apart from the root which has degree d2d-2; this will be relevant later for our path considerations.
We show a similar result for the size of the giant component for slightly non-regular degree sequences $\boldsymbol{d}$, where we allow a small set $S$ of vertices to have degree less than $d$; in our WSM setting, we will further be interested in the probability that a particular subset of the vertices $S^{\prime}$ does not belong to the giant. A variant of percolation we will use is the exact edge model for random graphs with a given degree sequence $\boldsymbol{d}=(d_{1},\dots,d_{n})$ and edge count $m\leq\tfrac{1}{2}\sum d_{i}$. Let $\tilde{\boldsymbol{G}}_{n,\mathbf{d},m}$ be the random graph model where each $i\in[n]$ is associated to $d_{i}$ half-edges; then we choose a subset of $2m$ half-edges and a matching of these uniformly at random; the final (multi)graph is obtained by projecting the edges on the vertex set $[n]$.
@@
    If $p<\tfrac{1}{d-1}$, $\mathbb{P}\big[{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{1/2}]\leq\mathrm{e}^{-\Omega(n)}$.
@@
    If $p>\frac{1}{d-1}$, $\mathbb{P}\big[{\mathcal{C}}_{1}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\leq(\chi-\delta)n]\leq\mathrm{e}^{-\Omega(n)}$ and $\mathbb{P}\big[{\mathcal{C}}_{2}(\tilde{\boldsymbol{G}}_{n,\boldsymbol{d},m})\geq Mn^{2/3}]\leq\mathrm{e}^{-\Omega(n)}$. Moreover, with $\varphi=\varphi(p)$ as in (5.1), for any subset $S^{\prime}\subseteq S$,
@@
Intuitively, the bound in the second item follows by revealing the $R$-neighbourhoods of vertices in $S^{\prime}$ for a suitably large constant $R$ (these are whp trees and disjoint from each other); then, for a vertex $i\in S^{\prime}$, the factor $(p\varphi+1-p)^{d_{i}}$ captures the probability that the percolation dies out within the neighourhood of the vertex $i$. The full proof of Lemma 5.1 is given in Appendix D.
Next, we consider the density of the edges inside the color classes. The following lemma says that in the ordered regime one colour class is supercritical, while the others subcritical. Recall that $\nu_{\mathrm{o}}$ and $\rho_{\mathrm{o}}$ are the vertex and edge statistics for the ordered phase respectively, defined in (2.1) and (2.5).
Lemma 5.2 ([14, Lemma 5.6]).
@@
Let $G=(V,E)$ be a graph. For a subset of edges $F\subseteq E$ and a vertex $v$, we let $F_{v}$ be the edges in $F$ incident to $v$. We will also consider the largest component $\mathcal{C}$} of the graph $(V,F)$, i.e. the component with the largest number of vertices (breaking ties arbitrarily). We say that $\mathcal{C}$} avoids a path $P$ if there is no vertex $v\in{\mathcal{C}}$ that belongs to $P$. The next lemma is the key technical step in our argument and will be used to show that, in the RC ordered phase, the largest component (which has size $\Omega(n)$ due to the supercriticality of colour 1, cf. Lemma 5.2) is “very close” to every vertex $v$.
@@
[…]
Then, for the RC ordered phase of $G$, $\pi_{G,\mathrm{o}}\big(\bigcup_{v\in V(G)}\mathcal{A}_{v}\big)\leq nd(d-1)^{\ell-1}(A+\varepsilon)^{\ell}$ where $A=\max_{i\in[q]}\big\{1-(1-{\sqrt{\hat{\varphi}_{1}})}\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}\big\}$ with $\rho_{\mathrm{o}},\nu_{\mathrm{o}}$ as in (2.1) and $\hat{\varphi}_{1}=\hat{\varphi}(p_{1})$ as in (5.1) for $p_{1}=p\frac{\rho_{\mathrm{o},11}}{\nu_{\mathrm{o},1}}$.
To parse the bound on $\pi_{G,\mathrm{o}}\big(\bigcup_{v\in V(G)}\mathcal{A}_{v}\big)$ note that the factor $nd(d-1)^{\ell-1}$ is just an upper bound on the total number of paths of length $\ell$. The factor $A^{\ell}$ controls the probability that a particular path $P$ avoids the largest component. The term $\hat{\varphi}_{1}$ in $A$ is the probability that the Galton-Watson process dies out from a vertex $v$ of the path that belongs to the supercritical color class 1 (note that $v$ has “remaining” degree d2d-2 other than its neighbours on the path, and that the relevant percolation parameter is given by $p_{1}$). The term $\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}$ accounts for the frequency that we encounter vertices in the supercritical color class 1 along the path $P$, which follows from the definition of the planting model. The exact quantitative dependence of $A$ to $\hat{\varphi}_{1}$ and $\frac{\rho_{\mathrm{o},i1}}{\nu_{\mathrm{o},i}}$ follows by a more technical calculation involving eigenvalues. Later, in Lemma 5.4, we will show that $A$ can be made less than $1/d^{5}$ by taking $q$ sufficiently large.
Proof of Lemma 5.3 (Sketch).
@@
From this, we obtain by Lemma 2.2 that $\mathbb{P}\big[\boldsymbol{F}\big(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big)\in\bigcup_{v\in V}\mathcal{A}_{v}\mid\boldsymbol{G}\big]\leq n\ln(n)d\big(d-1)^{\ell-1}(A+\tfrac{\varepsilon}{2}\big)^{\ell}$; since $\boldsymbol{F}\big(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}})$ has distribution $\hat{\pi}_{{\boldsymbol{G}},\mathrm{o}}$, Lemma 3.3 yields that, whp over $\boldsymbol{G}$, it holds that $\pi_{\boldsymbol{G},\mathrm{o}}(\bigcup_{v}\mathcal{A}_{v})\leq\hat{\pi}_{{\boldsymbol{G}},\mathrm{o}}(\bigcup_{v}\mathcal{A}_{v})+\mathrm{e}^{-\Omega(n)}\leq nd(d-1)^{\ell-1}(A+\varepsilon)^{\ell}$.
@@
Let $P$ be an arbitrary sequence of vertices $(v_{0},v_{1},\ldots,v_{\ell})$, together with a $q$-colouring $\tau=(\tau_{0},\ldots,\tau_{\ell})$. Let also $\xi$ be a set of paired half-edges that turn $v_{0},\ldots,v_{\ell}$ into a path in that order. Let $\boldsymbol{E}$ be the set of random edges in the graph before percolation. Denote by